# Not All Attention is Needed: Parameter and Computation Efficient Transfer Learning for Multi-modal Large Language Models

### v1

@@ Abstract
In this paper, we propose a novel parameter and computation efficient tuning method for Multi-modal Large Language Models (MLLMs), termed Efficient Attention Skipping (EAS). Concretely, we first reveal that multi-head attentions (MHAs), the main computational overhead of MLLMs, are often redundant to downstream tasks. Based on this observation, EAS evaluates the attention redundancy and skips the less important MHAs to speed up inference. Besides, we also propose a novel propagation-of-information adapter (PIA) to serve the attention skipping of EAS and keep parameter efficiency, which can be further re-parameterized into feed-forward networks (FFNs) for zero-extra latency. To validate EAS, we apply it to a recently proposed MLLM called LaVIN and a classic VL pre-trained model called METER, and conduct extensive experiments on a set of benchmarks. The experiments show that EAS not only retains high performance and parameter efficiency, but also greatly speeds up inference speed. For instance, LaVIN-EAS can obtain 89.98% accuracy on ScineceQA while speeding up inference by 2.2 times to LaVIN222The source code is anonymously released at: https://anonymous.4open.science/r/EAS-C0E5..
@@ 1 Introduction
Recently, the great success of Large Language Models (LLMs) also sparks an influx of interest in extending these giant model to more modalities, i.e., Multi-modal Large Language Models (MLLMs) (Luo et al., 2023b; Liu et al., 2023a; Li et al., 2023). However, compared with the unimodal LLMs, MLLMs inevitably involve much more computation with the introduction of new modalities (Bao et al., 2022; Alayrac et al., 2022; Li et al., 2023; Liu et al., 2023a), and their adaptions to downstream tasks also take more training and parameter expenditure (Kim et al., 2021b; Li et al., 2021, 2022). In this case, numerous efforts have been recently devoted to the efficient tuning of MLLMs (Zhang et al., 2023; Gao et al., 2023b; Wu et al., 2023b).
@@
Among these advancements, Wu et al. (2023b) propose a non-trivial task for large-scaled pre-trained models, termed parameter and computation efficient transfer learning (PCETL). A key intuition behind PCETL is that the sheer size of model parameters and computation is critical for large-scale pre-training but redundant to specific tasks. In this case, for downstream task adaptions, PCETL not only requires to reduce the number of updated parameters, akin to PETL (Houlsby et al., 2019; Hu et al., 2022; Sung et al., 2022b), but also needs to remove the redundant modules for better efficiency. Wu et al. (2023b) also propose a strong baseline called dynamic architecture skipping (DAS) to evaluate and skip the redundant Transformer layers with lightweight adapters, thereby saving both parameter and computation overhead on downstream tasks.
@@
Despite the effectiveness, we think that this strong baseline still leaves ample room to improve for MLLMs. Above all, DAS (Wu et al., 2023b) only considers the entire Transformer layer for redundancy evaluation, and a granular scheme deserves more explorations. As shown in Fig. 1-(a), multi-head attention (MHA) consumes 35% more time than feed-forward network (FFN), while its parameter size is only half. From existing literature (Martins et al., 2020; Tolstikhin et al., 2021), we also notice that MHA is mainly in charge of dependency modeling, and FFN is often used to improve model capacity. As shown in Fig. 1-(b), skipping a certain number of MHAs does not affect performance, while it will results in a drastic drop with FFN. In this case, we assume that FFN stores more pre-training knowledge for downstream task adaption. In contrast, for a simple task, the repeated attention modeling may be of limited significance. Based on these observations, we make an argument in this paper: “Not all attention is needed for the parameter and computation efficient tuning of MLLMs.”
@@
Another shortcoming of DAS is its skip connections with adapters, which still incurs non-negligible latency during inference. As revealed in recent progresses (Luo et al., 2023b; Gao et al., 2023b), the additional branches added by adapters inevitably slow down inference even with very limited FLOPs (Luo et al., 2023b). The same as shown in Fig.1, the latency caused by adapters takes up to 37.9% inference time of LaVIN (Luo et al., 2023b). This drawback somewhat undermines DAS’s efficiency for MLLMs.
@@
To overcome these limitations, we propose a novel approach for PCETL of MLLMs in this paper, termed efficient attention skipping (EAS). The principle of EAS is to only skip the redundant attention modeling for better efficiency and performance. To avoid the latency caused by adapters, we also equip EAS with an innovative parameter efficient module called propagation-of-information adapter (PIA), which can not only replace the skipped MHAs for adaption, but also can be seamlessly re-parameterized into FFNs for zero extra cost during inference. With these innovative designs, EAS can help MLLMs achieve better PCETL on vision-language tasks, i.e., true model acceleration with a small number of trainable parameters.
@@
To validate EAS, we first apply it to a recent MLLM called LaVIN (Luo et al., 2023b), and conduct extensive experiments on the open-set multi-modal question answer benchmark, namely ScienceQA (Lu et al., 2022). To align DAS (Wu et al., 2023b), we also apply EAS to a representative VL pre-trained model called METER (Dou et al., 2022) on three VL benchmarks, i.e., VQA2.0 (Goyal et al., 2017a), NLVR (Suhr et al., 2019a) and Flickr30K (Plummer et al., 2017). The experimental results show that EAS can not only retain high performance and parameter efficiency against existing PETL and PCETL methods (Zhang et al., 2023; Luo et al., 2023b; Wu et al., 2023b), but also boost the inference speed on downstream tasks. For instance, EAS can improve the inference speed by to DAS (Wu et al., 2023b) and to the default LaVIN (Luo et al., 2023b) without performance degradations.
@@
We propose a novel parameter and computation efficient tuning method for MLLMs, termed efficient attention skipping (EAS), which retains the high performance of MLLMs and reduces both parameter and computation expenditures on downstream tasks.
@@ 2.1 Vision-language pre-trained models
Recent years have witnessed the rapid development of vision-language pre-trained (VLP) models (Li et al., 2019; Lu et al., 2019; Su et al., 2019; Cho et al., 2021; Wang et al., 2022) on various vision-language (VL) tasks (Goyal et al., 2017b; Suhr et al., 2019b; Lu et al., 2022). These VLP models often adopt pre-trained vision and language models as the encoders, and build a deep fusion branch for cross-modal interactions. The merits of this design are that vision-language alignment can be effectively modeled in the additional fusion branch based on the well learned features of two modalities (Lu et al., 2019; Tan & Bansal, 2019; Li et al., 2021; Dou et al., 2022; Bao et al., 2022).
With the advent of large language models (LLMs) (Zhang et al., 2022; Touvron et al., 2023; OpenAI, 2023), recent researches (Luo et al., 2023b; Liu et al., 2023a; Li et al., 2023) extend these LLMs to vision-language tasks, i.e., multi-modal large language models (MLLMs), where LLMs are directly used for language modeling and multi-modal fusion. More recently, there are attempts of parameter-efficiently transferring LLM to vision-language tasks (Goyal et al., 2017b; Lu et al., 2022). These methods insert lightweight modules like adapter (Zhang et al., 2023) for downstream adaption instead of fully tuning LLMs, which however will slow down inference (Hu et al., 2022). Meanwhile, the use of visual tokens greatly exacerbates the already high computation of MLLMs (Liu et al., 2023a; Zhang et al., 2023). In this paper, we focus on both parameter and computation efficient tuning for MLLMs.
@@ 2.2 Parameter and Computation Efficient Tuning
To handle the ever-increasing parameters of large-scale pre-training models, parameter-efficient transfer learning (PETL) is proposed to save the training and storage costs of LLMs (Brown et al., 2020; Li & Liang, 2021b; Hu et al., 2022). The principle of PETL is to transfer the pre-trained models to downstream tasks with only a small number of trainable parameters (Brown et al., 2020; Li & Liang, 2021b), and its methodologies can be roughly divided into three main categories, i.e. prompt tuning (Jia et al., 2022b; Zhou et al., 2022), adapter (Karimi Mahabadi et al., 2021; Gao et al., 2023a) and the re-parameterized methods (Hu et al., 2022; Luo et al., 2023a). Very recently, Wu et al. (2023b) propose a new learning task called parameter and computation efficient transfer learning (PCETL), which has an additional requirement for model inference based on the settings of PETL. And a strong baseline called Dynamic Architecture Skipping (DAS) is also proposed to evaluates the redundancy of Transformer layers by a reinforcement learning method and replaces the most redundant layers with adapters. Motivated by (Hu et al., 2022; Wu et al., 2023b), we also focus on PCETL for MLLMs in this paper. We propose a granular exploration for MLLMs, i.e., the inner MHAs, and address the latency caused by using adapters for skip connections.
PCETL is also related but orthogonal to the network compression research (Lee et al., 2023; Shi et al., 2023; Liu et al., 2023b; Kim et al., 2021a). Similar with model quantization (Lee et al., 2023; Xiao et al., 2023; Wu et al., 2023c) and pruning (Shi et al., 2023; Nova et al., 2023; Jaiswal et al., 2023), PCETL aims to improve the model efficiency on downstream tasks. A main difference is that PCETL achieves this target companied with the PETL methods like adapter, and the original parameters of MLLMs should remain unchanged due to the expensive large-scale pre-training. In this case, the contributions of PCETL and network compression are orthogonal and can be further combined. For instance, the compact model obtained by the EAS method is full-precision and can be further accelerated using quantization methods (Dettmers et al., 2023).
@@ 3 Preliminary
We first recap the principle of parameter and computation efficient transfer learning (PCETL) (Wu et al., 2023b) for multi-modal large language models (MLLMs). Concretely, given an MLLM , which consists Transformer layers and its whole parameters denoted as , the objective of PCETL can be defined by
@@
From Eq.1, we can see that PCETL needs to reduce parameter cost during adaption, just like PETL (Li & Liang, 2021b; Hu et al., 2022; Sung et al., 2022b). Meanwhile, it also requires practitioners to evaluate and remove the redundant modules to speed up inference. Compared with previous model compression and acceleration tasks (Krishnamoorthi, 2018; Meng et al., 2020; Dong et al., 2021), one difference of PCETL is that the original model weights should keep unchanged, so as to facilitate more task adaptions, which is essential for the giant models that are expensively pre-trained with massive data.
Wu et al. (2023b) also propose a strong baseline called dynamic architecture skipping (DAS), which considers the entire Transformer layer for redundancy evaluation. As discussed above, the components in MLLMs often serve different roles, i.e., MHA and FFN. In this case, more detailed evaluations are beneficial for PCETL. Besides, the skip connections in DAS are aided by adapters (Houlsby et al., 2019; Sung et al., 2022b) which also incur non-negligible latency during the inference.
4 Efficient Attention Skipping
@@ 4.1 Overview
To achieve parameter and computation efficient tuning for MLLMs, we propose an Efficient Attention Skipping (EAS) approach in this paper, as illustrated in Fig. 3.
Concretely, given an MLLMs , EAS aims to remove its redundant components for downstream task adaption. Different to DAS (Wu et al., 2023b), the evaluation target of EAS is more specific, i.e., MHAs of a MLLM. Thus, the objective can be defined by
where represent a subset of MHA modules and denote all FFN modules. In practice, similar to DAS (Wu et al., 2023b), we first conduct a reinforcement learning based redundancy evaluation on MHAs for downstream tasks. Afterwards, given the redundancy scores, we determine which MHAs to skip, and replace them with an adapter-based connection.
However, adapters account for excessive inference time, as shown in Fig. 1-(a). And the widely used re-parameterized methods like LoRA (Hu et al., 2022) can approximate the QKV projections but cannot replace the entire MHA for skip connections.
@@ 4.2 Propagation-of-Information Adapter
When the redundant MHAs are evaluated, the next key step is how to effectively skip them without additional latency. To approach this target, we propose a novel re-paremeterizable adapter for EAS, termed Propagation-of-Information Adapter (PIA).
As shown in Fig. 3-(a), PIA also adopts a bottleneck structure to scale down the hidden dimension of the inputs, akin to existing adapters (Sung et al., 2022b; Luo et al., 2023a), thereby achieving a low-rank approximation of full tuning (Hu et al., 2022; Luo et al., 2023a; Wu et al., 2023a). However, it also differs in its inner path designs.
Concretely, given the input features of the layer, denoted as . PIA first projects onto two separate low-dimensional semantic spaces, and then obtain the hidden features of PIA by
@@
In addition to the low-dimension projection, Eq. 4 also realizes the information exchange among all input tokens via the combination with the averaged feature.
@@
Here, and are two linear projections, and is the path routing function defined by
where , is the weight matrix, is the temperature of softmax. This routing design stems from LaVIN (Luo et al., 2023b), and can handle different types of inputs, i.e., language-only or language-vision data.
Overall, the multi-path design can still hold parameter efficiency via the low-dimensional projections, while acting the role of MHA in information exchange.
Re-parameterization. With the effective structure, PIA also needs to re-parameterize its weights into the models to further improve computation efficiency. Here, we embed PIA into FFN, and avoid the skip connections (Wu et al., 2023b) to retain network complexity.
@@
To re-parameterize PIA into the nearby projection weights of FFN (Hu et al., 2022; Luo et al., 2022), we still need to convert PIA into one linear layer:
More specifically, we aim to simplify the definition of PIA in Eq. 4 and 5 into a two-layer linear MLP:
where and are the new projection matrices for down- and up-sampling, respectively, and , are biases.
@@
Similar to the down-sampling, the up-sampling, i.e. Eq.5, is redefined by
@@
Here, is the approximated low-rank weight matrix, and it then can be re-parameterized into the nearby weight matrices of FFN, like (Hu et al., 2022; Luo et al., 2023a).
Note that, the average feature in Eq. 4, i.e. , and the path router are example-dependent. In this case, the re-parameterization of PIA will be executed after the first decoding step of MLLMs. Considering the fact that most MLLMs often needs to decode a long sequence (Luo et al., 2023b; Zhang et al., 2023), PIA can still save massive computation during inference.
| NAT | SOC | LAN | TXT | IMG | NO | G1-6 | G7-12 | ||||
| Fully tuned methods | |||||||||||
| LLaVA-7B (Liu et al., 2023a) | 100.00% | 2.19s | 90.36 | 95.95 | 88.00 | 89.49 | 88.00 | 90.66 | 90.93 | 90.90 | 90.92 |
| Parameter-efficient methods | |||||||||||
| LLaMA-Adapter (Zhang et al., 2023) | 0.03% | 4.35s | 84.37 | 88.30 | 84.36 | 83.72 | 80.32 | 86.90 | 85.83 | 84.05 | 85.19 |
| LaVIN-7B (Luo et al., 2023b) | 0.06% | 3.70s | 89.25 | 94.94 | 85.24 | 88.51 | 87.46 | 88.08 | 90.16 | 88.07 | 89.41 |
| LaVIN-7B-LoRA (Hu et al., 2022) | 0.08% | 2.22s | 84.06 | 77.95 | 85.73 | 82.84 | 74.12 | 88.64 | 83.77 | 82.20 | 83.21 |
| Parameter and computation efficient methods | |||||||||||
| DAS-7B (Wu et al., 2023b) | 0.07% | 3.23s | 88.99 | 94.60 | 85.09 | 87.88 | 86.51 | 88.36 | 89.72 | 88.13 | 89.15 |
| EAS-7B (Ours) | 0.07% | 2.13s | 89.48 | 93.93 | 87.00 | 88.32 | 87.36 | 89.76 | 90.20 | 88.99 | 89.77 |
| EAS-7B (Ours) | 0.11% | 1.69s | 89.52 | 95.61 | 86.36 | 88.71 | 88.25 | 88.85 | 91.26 | 87.67 | 89.98 |

### v2

# Not All Attention is Needed: Parameter and Computation Efficient Tuning for Multi-modal Large Language Models via Effective Attention Skipping

@@ Abstract
Recently, Multi-modal Large Language Models (MLLMs) have garnered an influx of interest from both academia and industry. However, for the downstream task applications, MLLMs not only require to update a large number of parameters but also consume excessive computation. In this paper, we propose a novel parameter and computation efficient tuning method for MLLMs, termed Effective Attention Skipping (EAS). Concretely, we first reveal that multi-head attentions (MHAs) in MLLMs, the primary source of computation, are often redundant to downstream tasks. Based on this observation, EAS evaluates attention redundancy and skips the less important MHAs to speed up inference. Besides, we also propose a novel propagation-of-information adapter (PIA) to serve the attention skipping while maintaining parameter efficiency. More importantly, PIA can be further re-parameterized into feed-forward networks (FFNs) for zero-extra latency. To validate EAS, we apply it to a recently proposed MLLM called LaVIN, and conduct extensive experiments on a vision-language benchmark, namely ScienceQA. The experimental results show that EAS can not only retain the high performance of LaVIN but also reduce the updated parameters scale greatly while speeding up the inference speed to a large extent. For instance, LaVIN-EAS can obtain 89.98% accuracy while accelerating the inference speed by 2.2 times. Our code is given in https://github.com/DoubtedSteam/EAS
@@ 1 Introduction
Recently, the great success of Large Language Models (LLMs) also sparks an influx of interest in extending these giant models to more modalities, i.e., Multi-modal Large Language Models (MLLMs) [38, 32, 25]. However, compared with the unimodal LLMs, MLLMs inevitably require much more computation with the introduction of new modalities [3, 1, 25, 32], and their adaptions to downstream tasks also take more training and parameter expenditure [22, 28, 27]. In this case, numerous efforts have been recently devoted to the efficient tuning of MLLMs [54, 12, 53].
@@
Among these advancements, Wu et al.[53] propose a non-trivial task for large-scaled pre-trained models, termed parameter and computation efficient transfer learning (PCETL). A key intuition behind PCETL is that the sheer size of model parameters and computation is critical for large-scale pre-training but redundant to specific tasks. In this case, for downstream task adaptions, PCETL not only requires to reduce the number of updated parameters, akin to PETL [16, 17, 47], but also needs to remove the redundant modules for better efficiency. Wu et al.[53] also propose a strong baseline called dynamic architecture skipping (DAS) to evaluate and skip the redundant Transformer layers with lightweight adapters, saving both parameter and computation overhead on downstream tasks.
@@
Despite the effectiveness, we argue that this strong baseline still leaves ample room to improve. Above all, DAS [53] only considers the entire Transformer layer for redundancy evaluation, and a granular scheme deserves more explorations. As shown in Fig. 2-(a), multi-head attention (MHA) consumes 35% more time than feed-forward network (FFN), while its parameter size is only half. From existing literature [40, 49], we also notice that MHA is mainly in charge of dependency modeling, and FFN is often used to improve model capacity. As shown in Fig. 2-(b), skipping a certain number of MHAs does not affect performance, while it will results in a drastic drop when skipping FFNs. In this case, we think that FFN stores more pre-training knowledge for downstream task adaption. In contrast, for a simple task, the repeated attention modeling may be of limited significance. Based on these observations, we make an argument in this paper:
“Not all attention is needed for the parameter and computation efficient tuning of MLLMs.”
@@
Another shortcoming of DAS is its skip connections with adapters, which still incurs non-negligible latency during inference. As revealed in recent progresses [38, 12], the additional branches added by adapters inevitably slow down inference although they only brings some FLOPs[38]. The same as shown in Fig.2, the latency caused by adapters takes up to 37.9% inference time of LaVIN [38]. This drawback somewhat undermines DAS’s efficiency for MLLMs.
@@
To overcome these limitations, we propose a novel approach for the parameter and computation efficient tuning of MLLMs in this paper, termed Effective Attention Skipping (EAS). The principle of EAS is to only skip the redundant attention modeling for better efficiency and performance. To avoid the latency caused by adapters, we also equip EAS with an innovative parameter efficient module called propagation-of-information adapter (PIA), which can not only replace the skipped MHAs for adaption, but also can be seamlessly re-parameterized into FFNs for extra zero cost during inference. With these innovative designs, EAS can help MLLMs achieve better PCETL results on vision-language tasks, i.e., true model acceleration with a small number of trainable parameters.
@@
To validate EAS, we first apply it to a recent MLLM called LaVIN [38], and conduct extensive experiments on the open-set multi-modal question answer benchmark, namely ScienceQA [36]. To align DAS [53], we also apply EAS to a representative VL pre-trained model called METER [10] on three VL benchmarks, i.e., VQA2.0 [13], NLVR2 [45] and Flickr30K [43]. The experimental results show that EAS can not only retain high performance and parameter efficiency against existing PETL and PCETL methods [54, 38, 53], but also significantly speed up inference speed on downstream tasks. For instance, EAS can improve the inference speed by to DAS [53] and to the default LaVIN [38] without performance degradations.
@@
We propose a novel parameter and computation efficient tuning method for MLLMs, termed Effective Attention Skipping (EAS), which retains the high performance of MLLMs and reduces both parameter and computation expenditures on downstream tasks.
@@ 2.1 Vision-language pre-trained models
Recent years have witnessed the rapid development of vision-language pre-trained (VLP) models [29, 35, 44, 6, 51] on various vision-language tasks [14, 46, 36]. To facilitate vision-language alignment and learn generalized multi-modal representations, VLP models also adopt self-supervised objects and conduct pre-training on massive image-text pairs, such as masked language modeling [21, 34], masked image modeling [2, 5, 8] and image-text matching [10, 23].
With the advent of large language models (LLMs) [55, 50, 42], recent advance [38, 32, 25] focus on extending these LLMs to vision-language tasks, i.e., multi-modal large language models (MLLMs). For instance, BLIP-2[26] introduces a QFormer, to bridge the gap between vision and language modalities, selecting relevant visual semantics as additional tokens for LLMs. MINI-GPT4[57] uses a projection layer to map visual features into the language model. LLaVA [33] also follows this principle, but proposes a carefully designed training scheme in addition. Meanwhile, some efforts are also devoted to parameter-efficiently transfer learning of LLMs to vision-language tasks [14, 36]. These methods insert lightweight modules like adapter [54] for downstream adaption instead of fully tuning LLMs [17]. Meanwhile, the use of visual tokens greatly exacerbates the already high computation of MLLMs [32, 54]. In this paper, we focus on both parameter and computation efficient tuning for MLLMs.
@@ 2.2 Parameter and Computation Efficient Tuning
Parameter-efficient transfer learning (PETL) is proposed to save the training and storage costs of LLMs [4, 30, 17]. The principle of PETL is to transfer the pre-trained models to downstream tasks with only a small number of trainable parameters [4, 30], and its methodologies can be roughly divided into three main categories, i.e. prompt tuning [18, 56], adapter [20, 11] and the re-parameterized methods [17, 37]. Very recently, Wu et al.[53] propose a new learning task called parameter and computation efficient transfer learning (PCETL), which further requires to reduce model redundancy based on the setting of PETL. Wu et al.[53] also propose a strong baseline called Dynamic Architecture Skipping (DAS) for PCETL, which evaluates the redundancy of Transformer layers by a reinforcement learning method and replaces the most redundant layers with adapters [47]. Motivated by [17, 53], we also focus on PCETL for MLLMs in this paper. We propose a granular exploration for MLLMs, and address the latency caused by using adapters for skip connections.
@@ 3 Preliminary
We first recap the principle of parameter and computation efficient transfer learning (PCETL) [53] for multi-modal large language models (MLLMs). Concretely, given an MLLM , which consists Transformer layers, and its whole parameters denoted as , the objective of PCETL can be defined by
@@
From Eq.1, we can see that PCETL needs to reduce parameter cost during adaption, just like PETL [30, 17, 47]. Meanwhile, it also requires practitioners to speed up inference by evaluating and removing the redundant modules. Compared with previous model compression and acceleration tasks [24, 41, 7], one difference of PCETL is that the original model weights should keep unchanged, so as to facilitate more task adaptions, which is essential for the giant models that are expensively pre-trained with massive data.
Wu et al.[53] also propose a strong baseline called dynamic architecture skipping (DAS), which considers the entire Transformer layer for redundancy evaluation. As discussed above, the components in MLLMs often serve different roles, i.e., MHA and FFN. In this case, more detailed evaluations are beneficial for PCETL. Besides, the skip connections in DAS are aided by adapters [16, 47] which also incur non-negligible latency during the inference [53, 17].
4 Effective Attention Skipping
@@ 4.1 Overview
To achieve parameter and computation efficient tuning for MLLMs, we propose an Effective Attention Skipping (EAS) approach in this paper, as illustrated in Fig. 3.
Concretely, given an MLLMs , EAS aims to remove its redundant components for downstream task adaption. Different to DAS [53], the evaluation target of EAS is more specific, i.e., MHAs of a MLLM. Thus, the objective can be defined by
where represent a subset of MHA modules and denote all FFN modules. In practice, similar to DAS [53], we first conduct a reinforcement learning based redundancy evaluation on MHAs for downstream tasks. Afterwards, given the redundancy scores, we determine which MHAs to skip, and replace them with an adapter-based connection.
However, adapters will account for excessive inference time, as shown in Fig. 2-(a). The widely used re-parameterized methods like LoRA [17] can approximate the QKV projections but cannot replace the entire MHA for skip connections.
@@ 4.2 Propagation-of-Information Adapter
When the redundant MHAs are identified, the next key step is how to effectively skip them without additional latency. To approach this target, we propose a novel re-paremeterizable adapter for EAS, termed Propagation-of-Information Adapter (PIA).
As shown in Fig. 3-(a), PIA also adopts a bottleneck structure to scale down the hidden dimension of the inputs, akin to existing adapters [47, 37], thereby achieving a low-rank approximation of full tuning [17, 37, 52]. However, it also differs in its inner path designs.
Concretely, given the input features of the layer, denoted as , PIA first projects onto two separate low-dimensional semantic spaces, and obtain the hidden features of PIA by
@@
In addition to the low-dimension projection, Eq. 4 also realizes information exchange among all input tokens via the combination with the averaged feature.
@@
Here, and are two linear projections, and is the routing function defined by
where , is the weight matrix, is the temperature of softmax.
Overall, the multi-path design can still hold parameter efficiency via low-dimensional projections, acting the role of MHA in information exchange.
Re-parameterization. With the effective structure, PIA also needs to re-parameterize its weights into the models to further improve computation efficiency. Here, we embed PIA into FFN, and avoid the skip connections [53] to retain network complexity.
@@
To re-parameterize PIA into the nearby projection weights of FFN [17, 39], we still need to convert PIA into one linear layer:
More specifically, we aim to simplify the definition of PIA in Eq. 4 and 5 by
where and are new projection matrices for down- and up-sampling, and , are biases.
@@
Similar to down-sampling, up-sampling in Eq.5, is redefined by
@@
Here, is the approximated low-rank weight matrix, and it then can be re-parameterized into the nearby weight matrices of FFN, like [17, 37].
Note that, the average feature in Eq. 4, i.e. , and the path router are example-dependent. In this case, the re-parameterization of PIA will be executed after the first decoding step of MLLMs. Considering the fact that most MLLMs often needs to decode a long sequence [38, 54], PIA can still save massive computation during inference.